OpenStudy (anonymous):

how do you factor a^2+11k+18?

OpenStudy (jkristia):

11k or 11a ?

OpenStudy (anonymous):

If it is 11a then it would be (a+9)(a+2). a*a = a^2 9a+2a = 11a 9*2 = 18